Reports 1-1 of 1 Clear search Modify search
AdV-DET (Commissioning)
gouaty, berni - 15:29 Tuesday 21 November 2023 (62521) Print this report
OMC scan in single bounce

This morning, Francesco locked the OMC in single bounce and performed a temperature scan (Fig.1)

A zoom on Fig.2, shows that the power in the TEM00 was about 30 uW (read by B1_PD3), while the first order mode was at 5.9 uW and the second order mode at 0.45 uW.

Mode matching looks good, but we have a misalignment defect of the order of 20%. Offsets in the B5 QD2 quadrants were zeroed during this measurement.

Images attached to this report
Search Help
×

Warning

×